# Value-Penalized Auxiliary Control from Examples for Learning without Rewards or Demonstrations

### v1

@@ 6 Conclusion
In this work, we presented VPACE—value-penalized auxiliary control from examples, where we coupled scheduled auxiliary control with value penalization in the example-based setting to significantly improve learning efficiency and stability. Our experiments revealed that scheduled auxiliary control can exacerbate the learning of poorly-calibrated value estimates, which can significantly harm performance, and we alleviated this issue with an approach to value penalization based on the current value estimate of example data. We theoretically showed that our approach to value penalization still affords an optimal policy. We empirically showed that value penalization, together with scheduled auxiliary tasks, greatly improves learning from example states against a set of state-of-the-art baselines, including learning algorithms with other forms of feedback. Opportunities for future work include the further investigation of learned approaches to scheduling, as well as autonomously generating auxiliary task definitions.
Acknowledgments
We gratefully acknowledge the Digital Research Alliance of Canada and NVIDIA Inc., who provided the GPUs used in this work through their Resources for Research Groups Program and their Hardware Grant Program, respectively.
@@ Appendix A Reward Model Formulations
A.1 Learning a Reward Function
The most popular approach to reward modelling, known as inverse RL, tackles an unknown $R$ by explicitly learning a reward model. Modern approaches under the class of adversarial imitation learning (AIL) algorithm aim to learn both the reward function and the policy simultaneously. In AIL, the learned reward function, also known as the discriminator, aims to differentiate the occupancy measure between the state-action distributions induced by expert and the learner. In example-based control, the state-conditional discriminator loss is Eq. 3, where $D$ attempts to differentiate the occupancy measure between the state distributions induced by $\mathcal{B}^{*}$ and $\mathcal{B}$. The output of $D(s)$ is used to define $\hat{R}(s)$, which is then used for updating the Q-function using Eq. 6.
@@ A.1 Learning a Reward Function
A.2 Discriminator-Free Reward Labels with Mean Squared Error TD Updates
A simple alternative to using a discriminator as a reward model was initially introduced as soft-Q imitation learning (SQIL) in [6]. In standard AIL algorithms, $D$ is trained separately from $\pi$ and $Q^{\pi}$, where $D$ is trained using data from both $\mathcal{B}$ and $\mathcal{B}^{*}$, whereas $\pi$ and $Q^{\pi}$ are trained using data exclusively from $\mathcal{B}$. However, most off-policy algorithms do not require this choice, and approaches such as [4, 39] train $Q$, and possibly $\pi$, using data from both $\mathcal{B}$ and $\mathcal{B}^{*}$. It is unclear why this choice is often avoided in AIL, but it might be because it can introduce instability due to large discrepancy in magnitudes for Q targets given data from $\mathcal{B}$ and $\mathcal{B}^{*}$. Sampling from both buffers, we can define $\hat{R}(s_{t},a_{t})$ in Eq. 2 to be labels corresponding to whether the data is sampled from $\mathcal{B}$ or $\mathcal{B}^{*}$—in SQIL, the labels are respectively 0 and 1. The full training objective resembles the minimax objective in Eq. 3 where we set $D(s^{*})=1$ and $D(s)=0$.
@@ A.2 Discriminator-Free Reward Labels with Mean Squared Error TD Updates
A.3 Discriminator-Free Reward Labels with Binary Cross Entropy TD Updates
Eysenbach et al. [3] introduced recursive classification of examples (RCE), a method for learning from examples of success. RCE mostly follows the approach outlined above in Section A.2 but uses a weighted binary cross-entropy (BCE) loss with weights $1+\gamma w$ for data from $\mathcal{B}$ and $1-\gamma$ for data from $\mathcal{B}^{*}$. The TD targets are also changed from Eq. 11 and Eq. 12 to
@@ A.3 Discriminator-Free Reward Labels with Binary Cross Entropy TD Updates
A major benefit of this approach, compared with the MSE TD updates in Section A.2, is that $y(s,s^{\prime})\leq y(s^{*},s^{*})$ is always enforced at every update, meaning that our approach to value penalization would provide no extra benefit. Nonetheless, our results from Appendix B show that SQIL, even without value-penalization, almost always outperforms both RCE and SQIL with BCE loss (referred to SQIL-BCE here).
Appendix B Why Does SQIL Outperform RCE?
Eysenbach et al. [3] claimed that the highest performing baseline against RCE was SQIL, but it is worth noting that their implementation [^1] Available at https://github.com/google-research/google-research/tree/master/rce at time of writing. is a departure from the original SQIL implementation, which uses MSE for TD updates, described in Section A.2. Furthermore, Eysenbach et al. [3] also noted that it was necessary to add $n$-step returns to off-policy learning to get reasonable performance. In their experiments, however, this adjustment was not included for their version of SQIL with BCE loss. We complete the experiments using their implementation and show the average results across all RCE environments in Fig. 6, and also compare to using SQIL with MSE (without value penalization or auxiliary tasks) for TD updates.
These results empirically show that RCE and SQIL with BCE loss perform nearly identically, indicating that benefits of the changed TD targets and weights described in Section A.3 may not be as clear as previously described. Furthermore, SQIL with MSE clearly performs better on average, although it still performs worse than VPACE (see Appendix E). While these empirical results demonstrate that example-based control with BCE loss for TD updates is outperformed by SQIL with MSE, Eysenbach et al. [3] had theoretical results indicating that RCE should still learn an optimal policy. In the following section, we describe a flaw found in one of their proofs that may help to further understand why RCE is outperformed by SQIL with MSE.
B.1 Re-Examination of the Proofs from Section 4 of RCE [3]
In this section, we outline potential flaws in the proofs of Lemma 4.2 and Corollary 4.2.1 in RCE [3]. We then provide a lemma that shows RCE can be considered as recovering a specific reward function, thereby unifying RCE with other IRL algorithms.
To begin, recall that the RCE objective (i.e. Eq. (7) of Eysenbach et al. [3]) is defined as follows:
$\mathcal{L}^{\pi}(\theta):=p(e_{t+}=1)\mathbb{E}_{p(s_{t},a_{t}|e_{t+}=1)}\left[\log C^{\pi}_{\theta}(s_{t},a_{t})\right]+\mathbb{E}_{p(s_{t},a_{t})}\left[\log\left(1-C^{\pi}_{\theta}(s_{t},a_{t})\right)\right].$ (15)
Eysenbach et al. [3] stated the following Lemma and Corollary to demonstrate that optimizing Eq. 15 is equivalent to value iteration under the tabular setting:
Lemma 1.
(Lemma 4.2 of RCE) In the tabular setting, the expected updates for Eq. 15 are equivalent to performing value iteration with reward function $r(s_{t},a_{t})=(1-\gamma)p(e_{t}=1|s_{t})$ and a Q-function parameterized as $Q^{\pi}_{\theta}(s_{t},a_{t})=\frac{C^{\pi}_{\theta}(s_{t},a_{t})}{1-C^{\pi}_{\theta}(s_{t},a_{t})}$.
Corollary 1.
(Corollary 4.2.1 of RCE) RCE converges in the tabular setting.
In Lemma 1, we find that the proof intends to demonstrate equivalence with policy evaluation. In particular, the temporal difference (TD) target is defined to be $y=r(s,a)+\gamma\mathbb{E}_{s^{\prime},a^{\prime}}\left[Q(s^{\prime},a^{\prime})\right]$, which is optimizing the Bellman expected equation rather than Bellman optimal equation. Furthermore, the assignment equation for the ratio should in fact be an expectation over both the next state-and-action pairs (i.e. $s_{t+1}\sim p(\cdot|s_{t},a_{t})$ and $a_{t+1}\sim\pi(\cdot|s_{t+1})$) instead of only the next state, even if we aim to perform policy evaluation. As a result, Corollary 1 also fails to hold if we simply use the convergence proof of value iteration.
We now provide a lemma indicating that the optimal solution of Eq. 15 is equivalent to finding $Q^{\pi}$ with reward $r(s,a)=(1-\gamma)p^{\pi}(e=1|s)$:
Lemma 2.
Fix any policy $\pi:\mathcal{S}\to\Delta(\mathcal{A})$. Let $\theta^{*}=\operatorname*{arg\,min}_{\theta}\mathcal{L}^{\pi}(\theta)$ be the optimal solution of Eq. 15. Consider an MDP with reward function $r(s_{t},a_{t})=(1-\gamma)p(e_{t}=1|s_{t},a_{t})$. Then, in the tabular setting, $Q^{\pi}=\frac{C^{\pi}}{1-C^{\pi}}$ satisfies the Bellman expected equation. Furthermore, we have that $Q^{\pi}(s_{t},a_{t})=p^{\pi}(e_{t+}=1|s_{t},a_{t})$. That is, the solution of Eq. 15 is equivalent to finding $Q^{\pi}$.
Proof.
Fix any policy $\pi$. Consider an MDP with reward function $r(s,a)=(1-\gamma)p(e=1|s)$ and transition function $p(s^{\prime}|s,a)$. We define the Bellman expected operator $\mathcal{T}^{\pi}:\mathbb{R}^{\mathcal{S}\times\mathcal{A}}\to\mathbb{R}^{\mathcal{S}\times\mathcal{A}}$ to be
$\mathcal{T}^{\pi}q(s,a):=r(s,a)+\gamma\mathbb{E}_{s^{\prime}\sim p(\cdot|s,a),a^{\prime}\sim\pi(\cdot|s)}\left[q(s^{\prime},a^{\prime})\right].$
Note that under the tabular setting, $\mathcal{T}^{\pi}$ is a $\gamma$-contraction mapping under max-norm [40]. Thus, by Banach fixed-point theorem, we have that for any $q\in\mathbb{R}^{\mathcal{S}\times\mathcal{A}}$,
$Q^{\pi}=\left(\mathcal{T}^{\pi}\right)^{\infty}q.$
In other words, performing policy evaluation on $\pi$ with reward $r(s,a)=(1-\gamma)p(e=1|s)$ yields $Q^{\pi}$. Now, by Lemma 4.1 of Eysenbach et al. [3], we have that
$\frac{C^{\pi}(s_{t},a_{t})}{1-C^{\pi}(s_{t},a_{t})}=(1-\gamma)p(e_{t}=1|s_{t})+\gamma\mathbb{E}_{s_{t+1}\sim p(\cdot|s_{t},a_{t}),a_{t+1}\sim\pi(\cdot|s_{t})}\left[\frac{C^{\pi}(s_{t+1},a_{t+1})}{1-C^{\pi}(s_{t+1},a_{t+1})}\right].$ (16)
Eq. 16 satisfies the Bellman expected equation, thus by uniqueness of the fixed-point theorem, we have that
$Q^{\pi}(s_{t},a_{t})=\frac{C^{\pi}(s_{t},a_{t})}{1-C^{\pi}(s_{t},a_{t})}.$
Recall that $\frac{C^{\pi}(s_{t},a_{t})}{1-C^{\pi}(s_{t},a_{t})}$ is the Bayes’ optimal classifier for Eq. 15. Therefore, suppose we obtain the minimum of Eq. 15, that is, $\theta^{*}=\operatorname*{arg\,min}_{\theta}\mathcal{L}^{\pi}(\theta)$, then we have that
$p(e_{t+}=1|s_{t},a_{t})=\frac{C^{\pi}_{\theta^{*}}(s_{t},a_{t})}{1-C^{\pi}_{\theta^{*}}(s_{t},a_{t})}=\frac{C^{\pi}(s_{t},a_{t})}{1-C^{\pi}(s_{t},a_{t})}=Q^{\pi}(s_{t},a_{t}),$
meaning that finding the minimum of $\mathcal{L}^{\pi}(\theta)$ is equivalent to performing policy evaluation on $\pi$. ∎
With Lemma 2 and Lemma 4.3 of [3], we can view Algorithm 1 of Eysenbach et al. [3] as a form of policy iteration under the reward function $r(s,a)=(1-\gamma)p(e=1|s)$, where (i) updating the classifier $\theta$ corresponds to performing a policy evaluation step on $\pi$, and (ii) updating the policy $\pi$ corresponds to performing a policy improvement step. We note that, however, the RCE objective is not necessarily equivalent to policy evaluation when the true minimum of RCE, $\theta^{*}$, is not obtained.
Appendix C Finite-Sample Analysis of Value-Penalty Regularization
In this section, we show a finite-sample complexity bound of learning the Q-function via mean-squared error (MSE) with value-penalty regularization and demonstrate that we can use its solution for approximate policy iteration (API). Intuitively, given a fixed set of samples, a near-optimal solution obtained using the regularized loss is closed to the true optimal solution of the unregularized loss with high probability. Then, if we can control the error of the solution from the regularized loss, we can further bound the value error between the Q-function output by API and the optimal Q-function.
More formally, consider a parameterized function class $\mathcal{Q}=\{Q_{\theta}:\mathbb{R}^{\mathcal{S}\times\mathcal{A}}\to\mathbb{R}|\theta\in\Theta\}$. We define the MSE as
$\ell(\theta,Z_{N})=\frac{1}{N}\sum_{n=1}^{N}\left(Q_{\theta}(S_{n},A_{n})-G_{n}\right)^{2},$ (17)
where $Z_{N}=\{(S_{n},A_{n},G_{n})\}_{n=1}^{N}$ is $N$ samples of the state-action pairs and the corresponding expected return. The value-penalty regularizer is defined as
$h(\theta,Z_{n})=\lambda\frac{1}{N}\sum_{n=1}^{N}\left(\max(Q_{\theta}(S_{n},A_{n})-Q_{U},0)\right)^{2}+\left(\max(Q_{L}-Q_{\theta}(S_{n},A_{n}),0)\right)^{2},$ (18)
for some constants $Q_{L}<Q_{U}$, corresponding to the lowest and highest possible Q-values respectively, and positive constant $\lambda$. Then, the MSE with value-penalty regularization is defined as
$\ell_{reg}(\theta,Z_{N})$ $=\ell(\theta,Z_{N})+h(\theta,Z_{N}).$ (19)
During training, suppose we are given $N$ i.i.d. samples $Z_{N}$ from $\mathcal{D}$. We aim to approximate the regularized empirical risk minimizer (ERM) to obtain $\hat{\theta}$:
$\ell(\hat{\theta},Z_{n})+h(\hat{\theta},Z_{n})\leq\min_{\theta\in\Theta}\left[\ell(\theta,Z_{n})+h(\theta,Z_{n})\right]+\varepsilon^{\prime}.$ (20)
Our goal is to find the approximate regularized ERM $\hat{\theta}$ that achieves a low test loss, defined by
$\mathcal{L}(\theta,\mathcal{D})=\mathbb{E}_{Z\sim\mathcal{D}}\left[\ell(\theta,Z)\right].$ (21)
To begin, we assume that the function class has bounded outputs (or bounded parameters) which is a standard requirement for characterizing generalization errors:
Assumption 1.
(Bounded function output and realizability.) Let $C_{\mathcal{Q}}\geq\max(\lvert Q_{L}\rvert,\lvert Q_{U}\rvert)$. The parameterized function class $\mathcal{F}$ is bounded: $\mathcal{Q}=\{Q_{\theta}:\mathbb{R}^{\mathcal{S}\times\mathcal{A}}\to[-C_{\mathcal{Q}},C_{\mathcal{Q}}]|\theta\in\Theta\}$. Moreover, the optimal parameter of the test loss belongs to the function class: $\theta^{*}=\operatorname*{arg\,min}_{\theta}\mathcal{L}(\theta,\mathcal{D})\in\Theta$.
Note that the regularizer term Eq. 18 only has an effect when the function $Q_{\theta}$ predicts values that are out-of-range, therefore we set $C_{\mathcal{Q}}$ to be larger than the maximum possible returns for a more meaningful analysis, otherwise we recover the unregularized loss Eq. 17. Furthermore, we characterize the complexity of the function class $\mathcal{Q}$ using the expected Rademacher complexity with offset.
Definition 1.
The expected Rademacher complexity is defined to be
$\mathfrak{R}^{h}_{N}(\mathcal{Q},\mathcal{D})=\mathbb{E}_{Z_{N}\sim\mathcal{D}}\left[\mathbb{E}_{\sigma}\left[\sup_{\theta\in\Theta}\frac{1}{N}\sum_{n=1}^{N}\sigma_{n}\left(\ell(\theta,Z_{n})+0.5h(\theta,Z_{n})\right)-0.5h(\theta,Z_{N})\right]\right],$
where $\sigma_{1},\dots,\sigma_{N}$ are independent Rademacher random variables.
Then, by Corollary 6.21 of [41], we can show that the approximate ERM satisfies the oracle inequality with high probability, which we state here:
Lemma 3.
(Corollary 6.21 of [41].) Let $\Delta_{N}h(\theta)\leq\sup_{z,z^{\prime}}\left[h(\theta,z)-h(\theta,z^{\prime})\right]$. Assume that for some $C\geq 0$:
$\sup_{\theta\in\Theta}\left[\sup_{z,z^{\prime}}\left[\ell(\theta,z)-\ell(\theta,z^{\prime})+\Delta_{N}h(\theta)\right]\right]\leq C.$
Then the approximate ERM specified by Eq. 20, $\hat{\theta}$, satisfies the following oracle inequality. For any $\delta,\delta^{\prime}>0$, with probability $1-\delta-\delta^{\prime}$, we have that
$\mathcal{L}(\hat{\theta},\mathcal{D})\leq$ $\inf_{\theta\in\Theta}\left[\mathcal{L}(\theta,\mathcal{D})+\mathbb{E}_{Z_{n}}\left[h(\theta,Z_{N})\right]+\Delta_{N}h(\theta)\sqrt{\frac{\log(1/\delta^{\prime})}{2N}}\right]$
$+\varepsilon^{\prime}+2\mathfrak{R}^{h}_{N}(\mathcal{Q},\mathcal{D})+2C\sqrt{\frac{\log(1/\delta)}{2N}}.$
We note that the first term of RHS in Lemma 3 is zero due to 1. In particular, by 1 and uniqueness of Bellman expected operator, we have that $\mathcal{L}(\theta^{*},\mathcal{D})=0$. The second term $\mathbb{E}_{Z_{n}}\left[h(\theta,Z_{N})\right]=0$ and the first factor in the third term $\Delta_{N}h(\theta)=0$ since $Q_{\theta^{*}}$ is a valid Q-function, thus no clipping is applied. Without loss of generality, suppose that $\varepsilon^{\prime}=0$, then by Lemma 3 we have that with probability $1-\delta-\delta^{\prime}$,
$\mathcal{L}(\hat{\theta},\mathcal{D})\leq 2\mathfrak{R}^{h}_{N}(\mathcal{Q},\mathcal{D})+2C\sqrt{\frac{\log(1/\delta)}{2N}}=\lVert\varepsilon_{\text{approx}}\rVert_{2,\mathcal{D}}.$
In other words, as we increase the number of samples $N$, the approximation error of $\hat{\theta}$, $\lVert\varepsilon_{\text{approx}}\rVert_{2,\mathcal{D}}$, approaches zero with high probability. The boundedness in 1 further ensures that the Rademacher complexity is well-behaved and shrinks to zero as $N$ increases under the test loss [41].
Now, consider the approximate policy iteration algorithm for Q-function called AMPI-Q [42]. Instead of approximating the Q-function using the standard squared-loss Eq. 17, we instead use the regularized loss Eq. 19. From Lemma 3 we know that $\mathcal{L}(\hat{\theta},\mathcal{D})\leq\lVert\varepsilon_{\text{approx}}\rVert_{2,\mathcal{D}}$ with high probability. Consequently, at each iteration $k$ of AMPI-Q, we can replace the policy evaluation error $\lVert\varepsilon_{k}\rVert_{2,\mu}$ with $\lVert\varepsilon_{\text{approx}}\rVert_{2,\mathcal{D}}$ and immediately obtain a finite-sample analysis of AMPI-Q through Theorem 8 of [42].
Theorem 1.
With 1 and the assumptions from [42], after $k$ iterations, AMPI-Q with regularized loss Eq. 19 for policy evaluation satisfies the following with probability $1-\delta-\delta^{\prime}$:
$\lVert Q^{*}-Q^{\pi_{k}}\rVert_{2,\rho_{0}}\leq\frac{2(\gamma-\gamma^{k})(\mathcal{C}^{1,k,0}_{\infty})^{\frac{1}{2}}}{(1-\gamma)^{2}}\left(2\mathfrak{R}^{h}_{N}(\mathcal{Q},\mathcal{D})+2C\sqrt{\frac{\log(k/\delta)}{2N}}\right)+g(k),$
where $\mathcal{C}^{1,k,0}_{\infty}$ and $g(k)$ are the concentrability-coefficients related terms defined in [42].
Proof.
We can directly apply Theorem 8 of [42] (setting $p=2,q^{\prime}=1,q=\infty$) and take the union bound over the bad events, that is, with probability at most $(\delta+\delta^{\prime})/k,\mathcal{L}(\hat{\theta},\mathcal{D})>\lVert\varepsilon_{\text{approx}}\rVert_{2,\mathcal{D}}$ for each of $k$ iterations. ∎
Theorem 1 indicates that when we run AMPI-Q for sufficiently large number of iterations, the learned policy reaches optimal policy performance, even when using our proposed value-penalty regularization Eq. 18 on the standard squared loss for policy evaluation.
Appendix D Additional Environment, Algorithm, and Implementation Details
@@
D.1 Additional Environment Details
@@
D.2 Delta-Position Adroit Hand Environments
@@
D.3 Real World Environment Details
@@
Fig. 8 shows our experimental platform and setup for our two real world tasks. In both RealDrawer and RealDoor, the observation space contains the end-effector position, an ArUco tag [43] to provide drawer or door position (in the frame of the RGB camera; we do not perform extrinsic calibration between the robot and the camera), and the gripper finger position. The action space in both contains delta-positions and a binary gripper command for opening and closing. The action space for RealDrawer is one-dimensional (allowing motion in a line), while the action space for RealDoor is two-dimensional (allowing motion in a plane). The initial state distribution for RealDrawer allows for initializing the end-effector anywhere within a 10 cm line approximately 25 cm away from the drawer handle when closed. For RealDoor, the initial state distribution is a 20 cm $\times$ 20 cm square, approximately 20cm away from the door handle when closed. Actions are supplied at 5 Hz.
For both environments, for evaluation only, success is determined by whether the drawer or door is fully opened, as detected by the absolute position of the ArUco tag in the frame of the RGB camera. Our robot environment code is built on Polymetis [44], and uses the default hybrid impedance controller that comes with the library. To reduce environmental damage from excessive forces and torques, we reduced Cartesian translational stiffness in all dimensions from 750 N/m to 250 N/m, and the force and torque limits in all dimensions from 40 N and 40 Nm to 20 N and 20 Nm.
D.4 Additional Task Details
Fig. 9 shows representative images for all environments and tasks used in this work. Success examples for the Panda environments were gathered by taking $s_{T}$ from the existing datasets provided by [4]. Success examples for main tasks from the Sawyer environments were generated using the same code from [3], in which the success examples were generated manually given knowledge of the task. Auxiliary task data was generated with a similar approach. Success examples for the Adroit hand environments were generated from the original human datasets provided by [37]. Success examples for our real world tasks were generated by manually moving the robot to a small set of successful positions for each auxiliary task and main task.
As stated in Section 4.1, all Panda main tasks use the the auxiliary tasks release, reach, grasp, and lift.

### v2

@@ V Conclusion
In this work, we presented VPACE—value-penalized auxiliary control from examples, where we coupled scheduled auxiliary control with value penalization in the example-based setting to significantly improve learning efficiency. We showed that VPACE resolves Q overestimation, greatly improves the sample efficiency of example-based control against a wide set of baselines, and shows improved performance compared with with other forms of feedback. Opportunities for future work include the further investigation of learned approaches to scheduling, as well as autonomously generating auxiliary task definitions.
@@ Appendix A Reward Model Formulations
A-A Learning a Reward Function
The most popular approach to reward modelling, known as inverse RL, tackles an unknown $R$ by explicitly learning a reward model. Modern approaches under the class of adversarial imitation learning (AIL) algorithm aim to learn both the reward function and the policy simultaneously. In AIL, the learned reward function, also known as the discriminator, aims to differentiate the occupancy measure between the state-action distributions induced by expert and the learner. The learned reward function $\hat{R}$ is derived from the minimax objective [14, 16, 17]:
$\mathcal{L}(D)=\mathbb{E}_{\mathcal{B}}\left[\log\left(1-D(s)\right)\right]+\mathbb{E}_{\mathcal{B}^{*}}\left[\log\left(D(s^{*})\right)\right],$ (6)
where $D$ attempts to differentiate the occupancy measure between the state distributions induced by $\mathcal{B}^{*}$ and $\mathcal{B}$. The output of $D(s)$ is used to define $\hat{R}(s)$, which is then used for updating the Q-function.
@@ A-A Learning a Reward Function
A-B Discriminator-Free Reward Labels with Mean Squared Error TD Updates
[…]
@@ A-B Discriminator-Free Reward Labels with Mean Squared Error TD Updates
A-C Discriminator-Free Reward Labels with Binary Cross Entropy TD Updates
[3] introduced recursive classification of examples (RCE), a method for learning from examples of success. RCE mostly follows the approach outlined above in Section A-B but uses a weighted binary cross-entropy (BCE) loss with weights $1+\gamma w$ for data from $\mathcal{B}$ and $1-\gamma$ for data from $\mathcal{B}^{*}$. The TD targets are also changed from Eq. 7 and Eq. 8 to
@@ A-C Discriminator-Free Reward Labels with Binary Cross Entropy TD Updates
A major benefit of this approach, compared with the MSE TD updates in Section A-B, is that $y(s,s^{\prime})\leq y(s^{*},s^{*})$ is always enforced at every update, meaning that our approach to value penalization would provide no extra benefit. Nonetheless, our results from Appendix D show that SQIL, even without value-penalization, almost always outperforms both RCE and SQIL with BCE loss (referred to SQIL-BCE here).
Appendix B Additional Environment, Algorithm, and Implementation Details
@@
B-A Additional Environment Details
@@
B-B Delta-Position Adroit Hand Environments
@@
B-C Real World Environment Details
@@
Fig. 13 shows our experimental platform and setup for our two real world tasks. In both RealDrawer and RealDoor, the observation space contains the end-effector position, an ArUco tag [37] to provide drawer or door position (in the frame of the RGB camera; we do not perform extrinsic calibration between the robot and the camera), and the gripper finger position. Observations also include two stacked frames, in lieu of including velocity, to better follow the Markov property. The action space in both contains delta-positions and a binary gripper command for opening and closing. The action space for RealDrawer is one-dimensional (allowing motion in a line), while the action space for RealDoor is two-dimensional (allowing motion in a plane). The initial state distribution for RealDrawer allows for initializing the end-effector anywhere within a 10 cm line approximately 25 cm away from the drawer handle when closed. For RealDoor, the initial state distribution is a 20 cm $\times$ 20 cm square, approximately 20cm away from the door handle when closed. Actions are supplied at 5 Hz.
[…]
B-D Additional Task Details
Success examples for the Panda environments were gathered by taking $s_{T}$ from the existing datasets provided by [6]. Success examples for main tasks from the Sawyer environments were generated using the same code from [3], in which the success examples were generated manually given knowledge of the task. Auxiliary task data was generated with a similar approach. Success examples for the Adroit hand environments were generated from the original human datasets provided by [33]. Success examples for our real world tasks were generated by manually moving the robot to a small set of successful positions for each auxiliary task and main task. All Panda main tasks use the the auxiliary tasks release, reach, grasp, and lift.